y=|x2-2x-3| Solve the problem with the steps and draw the chart


1
Expert's answer
2020-12-29T14:00:47-0500
"y=|x^2-2x-3|"

"y=|(x+1)(x-3)|"

"y=0=>(x+1)(x-3)=0"

"=>x+1=0 \ \text {or}\ x-3=0"

"x_1=-1, x_2=3"

"x^2-2x-3\geq0" for "x\leq-1" or "x\geq3"


"x^2-2x-3<0" for "-1<x<3"


Therefore

"y = \begin{cases}
 x^2-2x-3 &\text{if } x\leq-1 \\
 -(x^2-2x-3) &\text{if } -1<x<3 \\
 x^2-2x-3 &\text{if } x\geq3
\end{cases}"